# SpikePingpong: High-Frequency Spike Vision-based Robot Learning for Precise Striking in Table Tennis Game

### v1

@@ 3.2 SONIC: Spike-Oriented Neural Improvement Calibrator
The SONIC methodology integrates ball trajectory data, velocity measurements, and physics-based predictions to precisely quantify the discrepancy between theoretical and actual hittable positions. For training purposes, we developed an extensive dataset that meticulously documents the systematic variations between physics-model predictions and empirically observed real-world interception positions, enabling our system to learn these complex error patterns.
Specifically, we first compute the required joint angles using inverse kinematics (IK) to move the center of the paddle to the hittable position predicted by the physics-based model. Subsequently, a high-speed Spike camera (dong2021spike, ) is employed to capture images of both the actual striking table tennis position and the paddle. Finally, the pixel distance between the ball’s position and the paddle center in these images is used to estimate the true spatial deviation. Concurrently, we record the ball’s position, velocity, and the hittable position predicted by the physics-based model.
@@
The SONIC network processes three key inputs: position and velocity vectors from the preceding $K$ frames, alongside the physics-based model’s predicted hittable position. Each input stream—position vectors, velocity vectors, and estimated hittable positions—is independently processed through dedicated multi-layer perceptrons (MLPs). These MLPs incorporate ReLU activation functions and strategic dropout regularization to enhance generalization capabilities. The extracted features are subsequently concatenated into a comprehensive representation, which is then refined through a Transformer encoder architecture to capture sophisticated temporal dependencies within the trajectory data. Finally, an output module composed of fully connected layers generates the predicted deviation vector. Figure 2 provides a detailed visualization of the SONIC model architecture.
@@
We optimize the model using the mean squared error (MSE) loss function, which minimizes the difference between the predicted and ground-truth deviation vectors.
$L_{MSE}(\theta)=\frac{1}{N}\sum_{i=1}^{N}||\hat{D}_{i}-D_{i}||^{2};\mathrm{where~{}}\hat{D}_{i}=f_{\theta}([p_{i},v_{i},h_{i}]).$ (3)
@@ 3.3 IMPACT: Imitation-based Motion Planning And Control Technology
After ensuring ball paddle interception through SONIC’s precise hittable position prediction, we propose IMPACT (Imitation-based Motion Planning And Control Technology) to derive the striking strategy for table tennis. This module learns strategic ball-hitting behaviors through imitation learning, enabling the robot to execute human-like tactical returns.
The IMPACT module requires accurate and diverse training samples that capture both the dynamic ball trajectories and ball landing region. Therefore, to collect such high-quality training data, we position the robotic arm at the hittable predicted by the SONIC using inverse kinematics (IK). Subsequently, based on the current state of the robotic arm moving to the hitting point, we add a randomly generated angle to three robot joints to execute the table tennis stroke. We record only those trials in which the ball is successfully returned on the opponent’s side of the table, and further classify them based on the specific landing region, enabling fine-grained control over hitting strategy.
The IMPACT module employs a transformer-based neural network that processes three types of input: ball trajectory data, robot joint configurations, and desired landing region signals (see Figure 2). Each modality is independently encoded into token representations using separate multi-layer perceptrons (MLPs). These tokens are then concatenated to form a unified input sequence, which is processed by a Transformer encoder to generate optimal adjustments for the joint angles.
@@
$L_{MSE}(\theta^{\prime})=\frac{1}{N}\sum_{i=1}^{N}||\hat{J}_{i}-J_{i}||^{2};\mathrm{where~{}}\hat{J}_{i}=f_{\theta^{\prime}}([p_{i},v_{i},j_{i},c_{i}]),j_{i}=\text{IK}(h_{i}).$ (4)
$f_{\theta^{\prime}}$ represents the transformer-based neural network with parameters $\theta^{\prime}$, $p_{i}\in\mathbb{R}^{K\times 3}$ denotes the position history, $v_{i}\in\mathbb{R}^{K\times 3}$ denotes the velocity history, $j_{i}\in\mathbb{R}^{6}$ represents the current 6-DOF robot joint configuration, $c_{i}\in\mathbb{R}^{4}$ represents the one-hot encoded control signal for the desired landing region, $J_{i}\in\mathbb{R}^{3}$ is the ground truth adjustment vector for the three critical joint angles, and $\hat{J}_{i}\in\mathbb{R}^{3}$ is the predicted joint adjustment vector.
Through this imitation-based approach, IMPACT enables the robot to adapt to different ball trajectories and target specific regions on the opponent’s court, transforming technical accuracy into tactical gameplay that significantly enhances the robot’s competitive capabilities in table tennis matches.
@@ 4 Experiments
In this section, we present comprehensive evaluations of SpikePingpong. In Section 4.1, we introduce our experimental setup, including dataset construction, implementation details, and evaluation metrics. Subsequently, we conduct extensive experiments to evaluate the ball-racket contact precision enabled by our SONIC module in Section 4.2. In Section 4.3, we assess the system’s spatial control capabilities for targeting across different table regions. Finally, we examine SpikePingpong’s ability to execute tactical sequences with consistent precision in Section 4.4, demonstrating the system’s reliability in complex gameplay scenarios.
@@ 4.1 Experiment Setting
We introduce the SpikePingpong Dataset for our table tennis system, containing two distinct parts. The first part includes 0.5k samples collected using a high-speed spike camera operating at 20 kHz, with each sample containing trajectory information (position and velocity vectors from K=3 consecutive frames), expected hittable position, and ground truth deviation vectors. The second part consists of 2k successful return demonstrations, each containing ball trajectory information, predicted hittable position converted to joint configurations via inverse kinematics, and the joint adjustments that resulted in effective returns to specific court regions. All data was collected using an automated ball launching machine generating shots with randomized trajectories, spins, speeds, and angles within the robotic arm’s operational range, under controlled lighting conditions with a standardized racket. This comprehensive dataset provides the foundation for training both our trajectory prediction correction model and strategic return planning system.
Figure 3: System Overview. Our system integrates three key subsystems: (1) a coordinate system for spatial tracking and transformation, (2) a multi-frequency control system with SONIC, IMPACT, and an EGM controller, and (3) a robot system based on the ABB IRB-120 arm equipped with a standard table tennis racket.
@@
As shown in Fig. 3, our system consists of three integrated subsystems. The coordinate system handles spatial transformations between world, camera, robot, and racket coordinates, with camera calibration performed before each session using ArUco markers to ensure accurate spatial measurements. The control system operates at multiple frequencies: the SONIC module processes ball trajectory data at 60Hz and predicts optimal interception positions, which are converted to joint configurations via inverse kinematics running at 20kHz. The IMPACT module operates at 2.4kHz to generate appropriate striking joint adjustments. These commands are transmitted to the EGM controller at 250Hz for responsive execution. The robot system employs an ABB IRB-120 robotic arm equipped with a standard table tennis racket, which receives and executes the control signals to perform the desired striking motions.
We implemented our comprehensive system on a workstation with an NVIDIA RTX 4090 GPU. For model training, the SONIC trajectory prediction module was trained for 2000 epochs using the Adam optimizer with a learning rate of 1e-3 and a cosine annealing learning rate schedule for optimal convergence. We used a batch size of 32, utilizing K=3 consecutive frames as input for robust trajectory prediction. The IMPACT strategic return module underwent similar rigorous training for 2000 epochs with identical optimization parameters but employed a reduced batch size of 4 to accommodate the complexity of the return planning task. All positional inputs were normalized to the range [0,1], while standard scaling was applied to velocity measurements to ensure numerical stability and consistent model performance across varying ball speeds and spin conditions.
@@
To evaluate our system’s precision and accuracy, we adopt quantitative metrics based on spatial accuracy following established methodologies (ding2022goalseye, ). For ball-racket contact precision, we measure the mean absolute error (MAE) and root mean square error (RMSE) for contact deviation prediction, analyzing both overall accuracy and axis-specific performance (Y-axis horizontal and Z-axis vertical predictions). Visual evidence from spike camera captures provides additional validation of interception precision.
Table 1: Ball Hittable Position Prediction Error. Our SONIC module achieves millimeter-level accuracy across both axes, demonstrating superior precision in predicting the actual ball-racket contact point.
Metric Value (mm)
Overall MAE 2.47
Overall RMSE 3.13
Y-axis MAE 2.98
Y-axis RMSE 3.69
Z-axis MAE 1.97
Z-axis RMSE 2.56
For landing accuracy, we divide the table tennis court into four specific regions, focusing on challenging aspects of ball placement control. We measure successful hit rate by calculating the percentage of balls landing within specified distances from the target position: primary accuracy zone (within 30cm of target center, 14% of reachable table area) and high precision zone (within 20cm of target center, 6% of reachable table area). This dual-threshold approach evaluates both general placement accuracy and high-precision targeting capabilities.
@@
Figure 4: Visualization of ball-racket contact precision with and without SONIC. Spike-based camera images show the ball center (green) and racket center (red) at contact moment, with deviation measurements in millimeters. The reduced offset in the with-SONIC condition demonstrates the module’s effectiveness in achieving more precise ball interception.
4.2 Ball-Racket Contact Precision Evaluation
Our SONIC module demonstrated significant improvements in predicting ball-racket contact deviations. Table 1 presents the quantitative results on the test set. The module achieved a mean absolute error (MAE) of 2.47mm and a root mean square error (RMSE) of 3.13mm for overall deviation prediction. Y-axis (horizontal) predictions showed superior accuracy with an MAE of 2.98mm, while Z-axis (vertical) predictions exhibited slightly lower error with an MAE of 1.97mm. As shown in Figure 4, SONIC substantially improves contact precision between ball and racket. The spike camera captures reveal a marked reduction in positional offsets when SONIC is activated, with minimal separation between ball center (green) and racket center (red), indicating highly accurate interception. This visual evidence corroborates the quantitative measurements and illustrates how SONIC’s millimeter-level prediction accuracy translates into improved physical contact precision.
4.3 Empirical Assessment of Spatial Control Capabilities
Figure 5: Spatial distribution of ball landing positions across four target regions. Each color represents a distinct target area, the concentric circles indicate the 20cm and 30cm precision thresholds.
We conducted rigorous comparative evaluations against established baseline methods to comprehensively validate our system’s precision capabilities across four strategically positioned target regions (A, B, C, and D). Table 2 presents a comprehensive analysis of success rates across various target regions and precision thresholds. SpikePingpong exhibits exceptional performance, achieving an average success rate of 91% within the 30cm accuracy zone and 71% within the more challenging 20cm high-precision zone, substantially surpassing all comparative methods. Our system convincingly outperforms both human average capabilities and established approaches including LFP (lynch2020learning, ) and GoalsEye (ding2022goalseye, ). Common vision-language-action models like ACT (zhao2023learning, ) show significantly lower accuracy when trained on our dataset. Diffusion Policy (chi2023diffusion, ) could not be evaluated due to its prohibitively high inference latency.
Table 2: Single-Target Return Accuracy: Success rates for ball placement across four distinct target regions (A-D) at both 30cm and 20cm precision thresholds. The table compares human players, previous robotic approaches, and our SpikePingpong system, with and without the SONIC prediction module. Higher percentages indicate better performance, with values presented as mean ± standard deviation where applicable.
Method A B C D Avg.
30cm (%) 20cm (%) 30cm (%) 20cm (%) 30cm (%) 20cm (%) 30cm (%) 20cm (%) 30cm (%) 20cm (%)
Human Avg. 26 8 30 10 32 18 30 16 30 13
LFP (lynch2020learning, ) 24 6 74 40 64 48 44 16 52 28
GoalsEye (ding2022goalseye, ) 58 30 75 58 44 32 36 14 53 34
Diffusion Policy (chi2023diffusion, ) 0 0 0 0 0 0 0 0 0 0
ACT (zhao2023learning, ) 11 ± 4 4 ± 2 12 ± 5 4 ± 1 10 ± 3 2 ± 1 14 ± 4 5 ± 2 12 ± 4 4 ± 2
SpikePingpong w/o SONIC 80 ± 2 57 ± 3 87 ± 3 65 ± 2 84 ± 2 46 ± 4 89 ± 1 72 ± 3 85 ± 3 60 ± 6
SpikePingpong 85 ± 3(27$\uparrow$) 65 ± 7(35$\uparrow$) 95 ± 5(20$\uparrow$) 71 ± 5(13$\uparrow$) 88 ± 2(24$\uparrow$) 66 ± 2(18$\uparrow$) 95 ± 1(51$\uparrow$) 83 ± 3(67$\uparrow$) 91 ± 5(38$\uparrow$) 71 ± 7(37$\uparrow$)
Table 3: Sequential Target Execution Performance: Success rates for 20-shot random target sequences. The Success Rate column shows the percentage of individual shots successfully reaching their designated targets within sequences, while subsequent columns present region-specific success rates at both 30cm and 20cm precision thresholds.
Method Success Rate(%) A B C D
30cm (%) 20cm (%) 30cm (%) 20cm (%) 30cm (%) 20cm (%) 30cm (%) 20cm (%)
Diffusion Policy (zhao2023learning, ) 0 0 0 0 0 0 0 0 0
ACT (zhao2023learning, ) 11 ± 3 7 2 11 5 9 5 14 6
SpikePingpong w/o SONIC 88 ± 4 80 65 92 58 76 65 90 78
SpikePingpong 95 ± 5(84$\uparrow$) 92(85$\uparrow$) 77(75$\uparrow$) 100(89$\uparrow$) 66(61$\uparrow$) 88(79$\uparrow$) 83(78$\uparrow$) 100(86$\uparrow$) 81(75$\uparrow$)
4.4 Sequential Target Execution Capabilities
To evaluate our system’s ability to execute tactical sequences, we designed a comprehensive experiment involving random target sequences spanning 20 consecutive returns (e.g., "BADCA…"). This challenging test rigorously assesses the system’s accuracy, reliability, and consistency throughout prolonged exchanges—a critical capability for realistic gameplay and competitive scenarios.
Table 3 presents the results of our multi-goal reaching experiment. SpikePingpong demonstrates exceptional performance, achieving an overall sequence success rate of 95% ± 5% across these 20-shot sequences. This means that approximately 95% of individual shots within these sequences successfully reached their designated targets. The substantial performance enhancement provided by SONIC becomes clear when examining the ablation results, as removing this component reduces the success rate to 88% ± 4%. These results demonstrate that our system not only excels at individual shot placement but can maintain precision while executing complex, extended tactical sequences. This capability represents a significant advancement toward robot table tennis systems that can implement sustained strategic gameplay rather than merely returning balls, suggesting potential for developing higher-level strategic planning modules.
4.5 Computational Efficiency
@@
Table 4: Computational Performance Comparison: Average inference times in milliseconds for generating return actions across different methods.
@@
Diffusion Policy (chi2023diffusion, ) 2437.22
ACT (zhao2023learning, ) 7.15
@@
Real-time responsiveness is paramount in table tennis robotics. As demonstrated in Table 4, our approach exhibits remarkable computational efficiency. SpikePingpong achieves an inference latency of merely 0.407ms, dramatically outpacing Diffusion Policy and substantially exceeding ACT’s performance. Our system’s rapid inference capability provides adequate time for physical actuation after decision-making, supporting responsive gameplay in real-world conditions.

### v2

@@ 3.1.2 System 2: Spike-Oriented Neural Improvement Calibrator
System 2 integrates ball trajectory data, velocity measurements, and physics-based predictions to precisely quantify the discrepancy between theoretical and actual hittable positions. For training purposes, we developed an extensive dataset that meticulously documents the systematic variations between physics-model predictions and empirically observed real-world interception positions, enabling our system to learn these complex error patterns.
Specifically, during each trial, we systematically record the ball’s 3D position and velocity vectors throughout its trajectory, along with the corresponding hittable position predicted by System 1’s physics-based model. Based on these predictions, we compute the required joint angles through inverse kinematics and execute the corresponding robotic arm motion to position the paddle center at the theoretically optimal hittable position. Subsequently, a Spike camera (Dong et al., 2021) captures images of the actual ball-paddle interaction at the moment of contact. The pixel distance between the ball’s observed position and the paddle center in these images provides a quantitative measure of the spatial deviation, which serves as the ground truth for System 2.
Network Architecture.
@@
System 2’s network processes three input modalities: historical position vectors $\mathbf{p}_{i}\in\mathbb{R}^{K\times 3}$, velocity vectors $\mathbf{v}_{i}\in\mathbb{R}^{K\times 3}$ from the preceding $K$ frames, and the physics-based predicted hittable position $\mathbf{h}_{i}\in\mathbb{R}^{3}$. Each modality is processed through dedicated MLPs with ReLU activation and dropout regularization for feature extraction. The concatenated features are then processed through a Transformer (Vaswani et al., 2017) encoder to capture temporal dependencies and contextual relationships across trajectory segments. Finally, a regression head with fully connected layers maps the refined representation to the predicted deviation vector.
Training Objective.
@@
We optimize the model using the mean squared error (MSE) loss function, which minimizes the difference between the predicted and ground-truth deviation vectors:
[…]
@@
Drawing inspiration from dual-system theory in cognitive science (Kahneman, 2011), this Fast-Slow system architecture combines the speed of immediate heuristic reasoning with the accuracy of experience-based learning, ensuring both real-time operation and precise task execution. Once trained, System 2 operates as a lightweight neural predictor that directly estimates deviation vectors from trajectory features without requiring spike camera feedback during deployment. This design enables the system to benefit from high-fidelity spike-based training data while maintaining computational efficiency and real-time performance in operational scenarios.
@@ 3.2 IMPACT: Imitation-based Motion Planning And Control Technology
Building upon our Fast-Slow system framework, where System 1 provides rapid physics-based predictions and System 2 delivers precise neural calibration through spike-oriented improvement, we introduce IMPACT (Imitation-based Motion Planning And Control Technology) for strategic striking behaviors. This module learns tactical ball-striking through imitation learning, enabling strategic returns that achieve targeted gameplay beyond mere interception.
Data Collection.
The IMPACT module requires high-quality training data that captures the complete striking process from ball trajectory to final landing outcomes. Firstly, we record the incoming ball trajectory and estimate the optimal hitting position using our Fast-Slow system framework. Based on this prediction, we compute the corresponding robot joint configurations through inverse kinematics and position the robotic arm accordingly. To generate diverse striking behaviors, we apply random angular perturbations to three critical robot joints before executing the stroke. We retain only successful trials where the ball returns to the opponent’s side, recording the perturbed joint angles and resulting landing positions. Each sample is labeled according to its specific landing region for fine-grained strategic control. This approach is highly efficient compared to teleoperation methods (Takada et al., 2022), as it leverages accurate hitting predictions to automatically position the robot, significantly reducing collection time while ensuring consistent data quality.
Network Architecture.
The IMPACT module employs a transformer-based neural network that processes three input modalities: ball trajectory sequences, robot joint configurations, and desired landing region specifications (see Figure 2). Each modality is independently encoded into token representations using dedicated multi-layer perceptrons (MLPs). These tokens are concatenated to form a unified input sequence, which is processed by a Transformer encoder that leverages self-attention mechanisms to capture inter-modal dependencies. The network outputs optimal joint angle adjustments for precise ball placement control, effectively integrating trajectory information, kinematic constraints, and strategic objectives within a unified framework.
Training Objective.
@@
[…]
Here, $f_{\theta^{\prime}}$ represents the neural network with parameters $\theta^{\prime}$, $p_{i}\in\mathbb{R}^{K\times 3}$ and $v_{i}\in\mathbb{R}^{K\times 3}$ denote the ball’s position and velocity history, $j_{i}\in\mathbb{R}^{6}$ represents the 6-DOF robot joint configuration, $c_{i}\in\mathbb{R}^{4}$ represents the one-hot encoded control signal for the desired landing region, $J_{i}\in\mathbb{R}^{3}$ is the ground truth adjustment vector, and $\hat{J}_{i}\in\mathbb{R}^{3}$ is the predicted joint adjustment vector.
Through this imitation learning framework, IMPACT enables the robot to dynamically adapt its striking strategy to varying ball trajectories while precisely targeting specific regions on the opponent’s court. This capability transforms the system from merely achieving technical accuracy to executing sophisticated tactical gameplay, significantly enhancing the robot’s strategic competitiveness in table tennis matches.
@@ 4 Experiments
We present comprehensive evaluations of SpikePingpong across multiple dimensions of table tennis performance. We begin with our experimental setup (Section 4.1), followed by the main results evaluating contact precision, tactical spatial control in both single and sequential tasks, and robustness under out-of-distribution and human-interaction scenarios (Section 4.2). Finally, we conduct ablation studies to validate our architectural choices (Section 4.3).
Figure 3: System Overview. Our system integrates three key subsystems: (1) a coordinate system for spatial tracking and transformation, (2) a multi-frequency control system with Fast-Slow system, IMPACT, and an EGM controller, and (3) a robot system based on the ABB IRB-120 arm equipped with a standard table tennis racket.
@@ 4.1 Experiment Setting
We introduce the training Dataset, collected using an automated launcher with highly randomized ball dynamics (spin, speed, and placement). It comprises two subsets: (1) 1k samples of trajectory and ground-truth contact deviation data, synchronously captured via RGB-D (60 Hz) and spike cameras (20 kHz); and (2) 2k expert return demonstrations, mapping ball trajectories to robot joint configurations for targeted placement. Further details are in Appendix D.
@@
As shown in Fig. 3, our system consists of three integrated subsystems. The coordinate system handles spatial transformations between world, camera, robot, and racket coordinates using ArUco marker calibration. The control system operates at multiple frequencies: Fast-Slow system processes trajectory data at 60Hz and converts predictions to joint configurations via inverse kinematics at 20kHz, while IMPACT operates at 2.4kHz for striking adjustments with commands transmitted to the EGM controller at 250Hz. The robot system employs an ABB IRB-120 robotic arm with a standard table tennis racket, implemented on a workstation with NVIDIA RTX 4090 GPU. Training details are provided in the Appendix E.
Baselines.
We benchmark against ACT (Zhao et al., 2023) and Diffusion Policy (Chi et al., 2023). To ensure fair comparison, we optimized both baselines by using state-based inputs (identical to IMPACT) instead of raw images to eliminate visual latency. Additionally, we accelerated Diffusion Policy using a 10-step DDIM sampler. All methods employ a synchronous just-in-time one-shot inference strategy, executing a single forward pass at the latest decision point to leverage the most accurate state estimation.
@@
We evaluate system performance through offline validation, online physical testing, and out-of-distribution (OOD) scenarios. Contact precision (MAE/RMSE) is measured on a held-out test set (80/10/10 split). Landing accuracy is assessed across four table regions (A, B, C, D) using two thresholds: 30cm (primary) and 20cm (high precision). Online evaluations utilize a ball launcher generating randomized trajectories (spin, speed, and placement) to ensure distinctness from the training set. To test generalization, we conduct OOD experiments with relocated launcher positions and zero-shot transfer against unseen human opponents. These standardized metrics ensure a rigorous assessment in the absence of publicly available hardware baselines.
@@
Figure 4: Visualization of ball-racket contact precision with and without System 2. Spike-based camera images show the ball center (green) and racket center (red) at contact moment. The reduced offset with the Fast-Slow System demonstrates improved ball interception accuracy.
Table 1: Ball Hittable Position Prediction Error. Our Fast-Slow system approach achieves superior precision in predicting the actual ball-racket contact point across both axes.
Method Y-axis Z-axis Overall
MAE RMSE MAE RMSE MAE RMSE
System 1 Only 53.65 60.39 34.62 38.45 44.13 50.62
RNN-based Method (Elman, 1990) 24.10 24.89 21.50 22.52 22.80 23.73
System 1 + System 2 (Ours) 9.87 11.16 14.82 16.10 12.34 13.85
4.2 Main results
System Precision and Efficiency
Accurate and timely ball interception is the prerequisite for any successful table tennis system. By leveraging our Fast-Slow system design, SpikePingpong achieves superior contact precision, with an MAE of 12.34 and an RMSE of 13.85 for deviation prediction (Table 1). As visualized in Figure 4, the spike camera captures reveal minimal spatial separation between the ball and racket centers at the moment of contact.
@@ 4.2 Main results
Table 2: Computational Performance Comparison. Average inference times in milliseconds for generating return actions across different methods.
@@
Diffusion Policy (Chi et al., 2023) 25.18
ACT (Zhao et al., 2023) 7.15
@@
Crucially, to translate this spatial accuracy into physical actuation, our system must operate with minimal latency. As shown in Table 2, SpikePingpong achieves an inference time of just 0.407ms, dramatically outpacing Diffusion Policy and ACT. This millisecond-level responsiveness ensures sufficient time for the robotic arm’s physical execution.
Table 3: Single-Target Return Accuracy (%). Success rates for ball striking across four distinct target regions (A-D) at both 30cm and 20cm precision thresholds. The table compares human players, previous robotic approaches, and our SpikePingpong system. Higher percentages indicate better performance. Standard deviations are denoted in subscripts.
Method Region A Region B Region C Region D Average
30cm 20cm 30cm 20cm 30cm 20cm 30cm 20cm 30cm 20cm
Human Avg. 48±6 28±4 52±5 32±3 56±7 38±5 54±4 34±6 53±6 33±5
Diffusion Policy (w/ vision) 3±2 1±1 4±3 0±0 2±2 1±1 3±2 0±0 3±2 1±1
Diffusion Policy (w/o vision) 6±3 2±1 7±4 2±2 5±2 1±1 6±3 1±1 6±3 2±1
ACT (w/ vision) 11±4 4±2 12±5 4±1 10±3 2±1 14±4 5±2 12±4 4±2
ACT (w/o vision) 18±5 7±3 20±6 8±3 17±4 6±2 19±5 7±2 19±5 7±3
SpikePingpong 91±3 69±4 93±2 72±3 92±4 70±5 93±3 71±4 92±3 70±4
Tactical Spatial Control
We conducted rigorous comparative evaluations against established baseline methods across four strategically positioned target regions (A, B, C, and D). Table 3 presents success rates across various target regions and precision thresholds. SpikePingpong exhibits exceptional performance, achieving an average success rate of 92% within the 30cm accuracy zone and 70% within the 20cm high-precision zone, with consistent performance across all target regions. Our system substantially outperforms the human average (53% at 30cm) and all baseline methods. Notably, while baselines like ACT (Zhao et al., 2023) performed better without visual inputs (from 12% to 19%), our approach still demonstrates a significant performance leap.
Beyond isolated returns, to evaluate our system’s ability to execute tactical sequences, we designed experiments involving random target sequences spanning 100 consecutive returns across four target regions, with 25 targets per region presented in random order. We define success using a 30cm precision threshold, with additional 20cm precision metrics provided for detailed analysis. Table 4 presents the results of our sequential target execution experiment. SpikePingpong achieves an overall sequence success rate of 78%, significantly outperforming the human baseline of 45%. Our system demonstrates balanced performance across all court regions with consistent precision under stricter 20cm thresholds. These results demonstrate that our system maintains high precision while executing complex, extended tactical sequences that surpass human-level performance, representing a significant advancement toward sustained strategic gameplay rather than merely returning balls.
Table 4: Sequential Target Execution Performance (%). Success rates for 100-shot random target sequences. The overall rate column shows the percentage of individual shots successfully reaching their designated targets (30cm), while preceding columns present region-specific success rates. Standard deviations for the overall rate are denoted in subscripts.
Method Region A Region B Region C Region D Success Rate
30cm 20cm 30cm 20cm 30cm 20cm 30cm 20cm
Human Avg. 44 26 47 28 43 25 46 29 45±5
Diffusion Policy (w/ vision) 1 0 2 0 1 0 1 0 1±1
Diffusion Policy (w/o vision) 2 1 3 1 2 0 2 0 2±2
ACT (w/ vision) 7 2 9 3 6 2 10 3 8±3
ACT (w/o vision) 14 5 16 6 13 4 17 5 15±4
SpikePingpong 76 52 79 54 77 51 80 55 78±3
Robustness and Out-of-Distribution Generalization
To evaluate the generalization capability of our policy beyond the training distribution, we conducted a challenging out-of-distribution (OOD) experiment. While all training data was collected with the ball launcher at a fixed central position on the opponent’s table, for the OOD test, we physically moved the launcher to two new, unseen off-center positions. This fundamentally altered the entire distribution of incoming ball trajectories, including their angles, speeds, and bounce locations.
Table 5: Out-of-Distribution Generalization. Success rates on seen vs. unseen trajectory distributions.
Condition Avg. Success Rate
30cm (%) 20cm (%)
In-Distribution 92 ± 3 70 ± 4
Out-of-Dist. 74 ± 5 52 ± 6
As shown in Table 5, while there is a predictable drop in performance compared to the in-distribution setting, our system maintained a remarkable average success rate of 74% for 30cm targets. This result is significant, as the off-center launcher positions create entirely novel trajectories that the policy has never encountered. The sustained high performance strongly suggests that our policy has learned a robust, underlying model of ball dynamics and striking control, rather than simply memorizing patterns from the training data. This validates the generalization capabilities of our framework.
Table 6: Adaptation and Generalization to Human Demonstrations (%). Success rates on precision targeting tasks. ”Seen Player” refers to testing on the same human demonstrator whose data was used for fine-tuning. ”Unseen Player” refers to zero-shot testing on a new human player. Standard deviations are denoted in subscripts.
Precision Condition Region A Region B Region C Region D Average
30cm Seen Player (Person A) 51±5 47±3 42±3 50±4 47±3
Unseen Player (Person B) 35±6 31±5 27±5 32±4 31±5
20cm Seen Player (Person A) 29±2 23±4 27±2 28±3 27±3
Unseen Player (Person B) 18±3 13±4 16±3 14±4 15±4
Furthermore, to push the boundaries of generalization, we tested our system’s ability to adapt to and generalize from complex, human-generated trajectories. Unlike the structured patterns from a robotic launcher, human shots introduce significant, unstructured variability, representing a more challenging data distribution. We conducted a two-stage experiment to assess this capability, with results summarized in Table 6. First, to test adaptability, we fine-tuned our model on a small dataset of 100 demonstrations from a single human player (Person A). As shown in the ”Seen Player” rows of Table 6, when tested on this same player, the system achieved a meaningful success rate of 47% for 30cm targets, demonstrating strong sample efficiency. Next, to perform a stricter test of generalization, we evaluated this fine-tuned model in a zero-shot setting against a new, unseen player (Person B). The results in the ”Unseen Player” rows show a drop in performance to 31%, as expected in a challenging OOD scenario. Nevertheless, achieving over 31% accuracy on a completely new person without any specific fine-tuning is a non-trivial result. It provides promising evidence that our model captures generalizable features of human-like dynamics rather than simply overfitting to an individual’s style, highlighting its potential for collaborative human-robot scenarios.
4.3 Ablation Study
We conducted ablation experiments across four target regions (A, B, C, D) within a 30cm radius threshold to evaluate each component’s contribution. Our Fast-Slow system architecture achieves 92% accuracy in single-target returns, representing a 25-percentage-point improvement over the RNN (Elman, 1990) baseline. For sequential target execution, our Fast-Slow system achieves 78% success rate compared to 52% for the RNN-based method (Elman, 1990). This substantial enhancement stems from the system’s ability to account for ball-racket interaction sensitivity, where identical striking motions can produce vastly different trajectories depending on precise contact positioning. The consistent improvements across both single-target and sequential tasks validate the robustness of our architectural design.
Table 7: Ablation Study (%). Performance comparison of different trajectory prediction components in our SpikePingpong system across two distinct tasks.
Method Region A Region B Region C Region D Overall
Task 1: Single-Target Return Accuracy
System 1 + IMPACT 22±6 25±5 21±7 24±5 23±6
RNN (Elman, 1990) + IMPACT 65±5 68±4 66±6 69±3 67±5
SpikePingpong 91±3 93±2 92±4 93±3 92±3
Task 2: Sequential Target Execution Performance
System 1 + IMPACT 18±5 20±4 17±6 19±5 15±4
RNN (Elman, 1990) + IMPACT 55±4 58±5 54±6 57±4 52±6
SpikePingpong 76±4 79±3 77±4 80±3 78±3